OpenStudy (a1234):

Can someone please explain how the answer is 143?

OpenStudy (zzr0ck3r):

This is very meta.

OpenStudy (amistre64):

h is 90 degrees, and we can trig the rest

OpenStudy (a1234):

How do we trig?

OpenStudy (amistre64):

by some thrm, arc(HK) is twice angleG

OpenStudy (amistre64):

we have a right triangle, and know leg lengths ....

OpenStudy (amistre64):

 trig to find \(\alpha\) and \(\beta \)

OpenStudy (amistre64):

what trig function uses just the legs?

OpenStudy (a1234):

tan

OpenStudy (a1234):

so tanx = 12/4

OpenStudy (amistre64):

then tan it tan(a) = 12/4 inverse for the angle

OpenStudy (a1234):

Got it! Thanks very much!

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

the required arc is therefore \(2\alpha\) :)